OpenStudy (jade27p):

What is the equation, in slope-intercept form, of the line that passes through (0, 6) and has a slope of 4?

Vocaloid (vocaloid):

Medals 0 the slope-intercept formula is y = mx + b to find b, replace "y" with your y-coordinate (6), replace "x" with your x-coordinate (0), replace "m" with your slope (4), and solve for b

OpenStudy (jade27p):

6x+4?

OpenStudy (jade27p):

=y

Vocaloid (vocaloid):

replace "y" with 6 replace "x" with 0 replace "m" with 4

OpenStudy (jade27p):

will they all stay positive

Vocaloid (vocaloid):

what?

OpenStudy (jade27p):

the numbers

Vocaloid (vocaloid):

all you are doing is taking y, x, and m and replacing them with these new numbers

Vocaloid (vocaloid):

erase the y and write 6 where it used to be erase the x and write 0 where it used to be erase the m and write 4 where it used to be

OpenStudy (jade27p):

y=6 x=0 m=4

Vocaloid (vocaloid):

no, in the equation

Vocaloid (vocaloid):

take the equation y = mx + b

Vocaloid (vocaloid):

replace y with 6 replace x with 0 replace m with 4

OpenStudy (jade27p):

6=(4,0)+b

Vocaloid (vocaloid):

almost

Vocaloid (vocaloid):

6 = 4*0 + b

Vocaloid (vocaloid):

do you know how to find what b equals?

OpenStudy (jade27p):

6?

Vocaloid (vocaloid):

good!

Vocaloid (vocaloid):

now we know that b = 6

Vocaloid (vocaloid):

so we go back to our original equation y = mx + b we know that m = 4 and b = 6 so our answer is y = 4x + 6

OpenStudy (jade27p):

y=4x+6

Vocaloid (vocaloid):

good!

OpenStudy (jade27p):

thank you, can you help with 1 more

Vocaloid (vocaloid):

sure

OpenStudy (jade27p):

Use the graph below to answer this question: graph of line going through negative 1, 5 and 4, negative 5 Find the average rate of change for the given function from x = −1 to x = 4.

Vocaloid (vocaloid):

use the formula f(4) - f(-1) -------- 4 - (-1)

Vocaloid (vocaloid):

f(4) = -5 f(-1) = 5

Vocaloid (vocaloid):

so your formula becomes 5 - (-5) ------ 4 - (-1) all you need to do is simplify that and you'll have your answer

Vocaloid (vocaloid):

(------) means divide

OpenStudy (jade27p):

hold on one sec i need to write this

OpenStudy (jade27p):

i got 2

Vocaloid (vocaloid):

oh whoops I made a typo

Vocaloid (vocaloid):

-5 -(5) ------- 4 - (-1) so it ends up being -2

OpenStudy (jade27p):

ohhh okay, thank you